OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 6x – 2y = 2 –3x + 4y = 5 (1, 2) (1, 3) (2, 3) (1, –2)

OpenStudy (hhmorris98):

(1,2)

OpenStudy (michele_laino):

Hint: I multiply the first equation by \(2\), so I get this: \[\left\{ \begin{gathered} 12x - 4y = 4 \hfill \\ \hfill \\ - 3x + 4y = 5 \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

Now, please add such equations together, what do you get?

OpenStudy (anonymous):

i still don't get it

OpenStudy (michele_laino):

If I add such equations together, I can write this: \[12x - 4y - 3x + 4y = 4 + 5\] and after a simplification, I get: \(9x=9\) next I divide both sides by \(9\): \[\frac{{9x}}{9} = \frac{9}{9}\] please simplify

OpenStudy (anonymous):

its 1

OpenStudy (michele_laino):

that's right! we get \(x=1\)

OpenStudy (michele_laino):

now, if I replace \(x=1\) into your first original equation, I get: \[6 \cdot \left( 1 \right) - 2y = 2\] please solve for \(y\)

OpenStudy (michele_laino):

hint: after a simple computation, I get: \(6-2y=2\)

OpenStudy (anonymous):

its 4

OpenStudy (michele_laino):

I'm sorry, it is not \(y=4\) Hint: next I add \(-6\) to both sides, so I get: \[6 - 2y - 6 = 2 - 6\]

OpenStudy (michele_laino):

after a simplification I can write: \(-2y=-4\) then I divide both sides by \(-2\): \[\frac{{ - 2y}}{2} = \frac{{ - 4}}{2}\] please simplify

OpenStudy (anonymous):

2

OpenStudy (michele_laino):

oops.. I meant: \[\frac{{ - 2y}}{{ - 2}} = \frac{{ - 4}}{{ - 2}}\] That's right! we have \(y=2\)

OpenStudy (michele_laino):

so, what is the right option?

OpenStudy (michele_laino):

please keep in mind that the solution is given by the ordered pair \((1,2)\)